4 m 4 m 4 m Ey Ex E F H 2.5 m D A B С 3 kN 3 kN 3 kN

All forces pls
Answer #1

ol um Eye um H Exe F 10. 2.5m B D Toy 3 KN 3KN 3 KN het To calculate forces in all member, calculate the suppost reaction - EJoint-D coso = 0.848, sino = 0.53 tano = 2.5 Efx=0 & FDG Coso =0 FoG=0 FOG FDH 1 3 KN sfy=0 3 FDH & Forsino +3=0 → FOH = -3KN> Flor= - 3 KN (compressive) Joint be = 0 FGF GY FGH - FEG + FBG sino =0 o FBG Fo G=0 - FB G X 0.53 = 3 FBG= 5.66 KN FCol (Te→ z Fx = 0 - FBG coso = FBE Coso FBE- 5.66 KN (Tensile) Sty=0 FBF + FBG sino & FBE sino -3=0 FBFT 5.668 0.53+ 5.668 0.53-3=0

Member Force(KN) Nature of force
DG 0 Neutral
DH 3 COMPRESSIVE
CH 5.66 TENSILE
GH 4.8 COMPRESSIVE
CF 5.66 TENSILE
CG 3 COMPRESSIVE
BG 5.66 TENSILE
GF 9.6 COMPRESSIVE
BE 5.66 TENSILE
BF 3 COMPRESSIVE
AF 0 Neutral
EF 4.8 COMPRESSIVE
AE 3 TENSILE